THE EFFECTIVE HAMILTONIAN FOR THE GROUND STATE OF ^{207}P b^{19}F AND NEW MEASURE-MENTS OF THE FINE STRUCTURE SPECTRUM NEAR 1.2 μm.

RICHARD MAWHORTER, BENJAMIN MURPHEY, ALEXANDER BAUM, Department of Physics and Astronomy, Pomona College, Claremont, CA 91711; \underline{TREVOR J. SEARS}, Chemistry Department Brookhaven National Laboratory, Upton, NY 11973 and Stony Brook University, Stony Brook, NY 11794; T. ZH. YANG, P. M. RUPASINGHE, C. P. MCRAVEN^{a}, N. E. SHAFER-RAY, Homer L. Dodge Department of Physics and Astronomy, University of Oklahoma, Norman, OK; LUKAS D. ALPHEI AND JENS-UWE. GRABOW, Gottfreid-Wilhelm-Liebniz-Universität, Institut für Physikalische Chemie & Elektrochemie, D-30167 Hannover, Germany.

We have measured rotational transitions in the ground, X_{1} ^{2}Π_{1/2}, electronic state of naturally occuring isotopomers of PbF in a supersonic free jet Fourier transform microwave spectrometer. The data for ^{207}P b^{19}F is particularly interesting because it is a candidate for a future experimental e-EDM measurement. To fit the data for this species to the measurement precision, the nuclear spin-spin dipolar interaction and a second term that can be equivalently viewed as a centrifugal distortion correction to the familiar Frosch and Foley hyperfine coupling terms, or an Ω− dependent correction to the nuclear spin-rotational coupling are required, in addition to the standard terms. To characterize the higher X_{2} ^{2}Π_{3/2} component of the ground state of PbF, we are attempting a direct measurement of transitions between the two components in a slit jet-cooled sample using a frequency comb-referenced extended cavity diode laser. This spectrum was originally detected in a hot source by Fourier transform near-infrared spectroscopy,^{b} but low−J transitions were unresolved at that time. Acknowledgments: Work at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy. Work by N. E. Shafer-Ray was performed with support from the National Science Foundatation award NSF-0855431. J.-U. Grabow ackonwledges funding from the Deutsche Forschungsgemeinschaft and the Land Niedersachsen.

aCurrent Address: Chemistry Department, Brookhaven National Laboratory, Upton, NY 11973

bK. Ziebarth, K. Setzer, O. Shestakov and E. Fink J. Molec. Spectrosc. \underline{191}, 108 1998.